OpenStudy (anonymous):

Find dF/dC. C=5/9(F-32).

OpenStudy (anonymous):

Just solve for F and then differentiate with respect to C. \[\frac{9}{5}C+32=F \implies \frac{dF}{dC}=\frac{9}{5}\]

OpenStudy (anonymous):

That's what I have, but the book says the answer is -5, 3...